A working party of Rob, Graham, Ben and TRD at BY today. Rob started to check oil levels in all gear boxes, some needed topping up but generally they had not used much ATF or F/dive oil since the last check about two months ago. Graham spent time sorting out land line charging cables which were in a tangle. He has rewired the supplies so that both trailer car batteries can be charged together. He watered the tanks ready for this evenings SECR. Ben arrived later then he and Rob dealt with an air leak on the wiper control second man side of 51941 cab. The supply pipe had broken at the control inlet. They selected a new control from our Container and fitted it in position. They exchanged the wiper blades for new ones. Then tried both wipers on air pressure to check speed and blade sweep. Later we all did some tidying up moving non-essential items to container to create room in our small shed, swept the floor and went for lunch. We finished early because most of us had afternoon appointments to go to. For the short amount of time we were on site a lot of useful things were achieved.